how much energy is emitted out when amole of electron falls in 1g


atom of H atom II to I orbit?


Rh=109678 cm-1

Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Yash Jain,

Ans:- We know that

1/λ = Rh(1/(n1)² - 1/(n2)²)

putting n1=1 n2=2 and Rh as given, we get the value of λ. Then find energy putting E=hc/λ

Then multiply it with the Avogadro number and get the total energy

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Yash Jain !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur